OpenStudy (anonymous):

the cruise control of megan's truck is set at 65 miles per hour, but her actual speed may vary by as much as 5 miles per hour. write an inequality that can be used to determine the actual speeds at which megan could be traveling

OpenStudy (anonymous):

\[ |v - 65| \leq 5 \]

OpenStudy (abb0t):

I hope there's no cops around megan, she might get a ticket driving at such high speeds.

OpenStudy (anonymous):

thanks @wio

OpenStudy (anonymous):

@brokenbells83 Do you understand the inequality?

OpenStudy (anonymous):

@wio to be honest ... not really

OpenStudy (anonymous):

Basically, if you have two values, \(v\) her actual speed and \(65\) her typical speed, and you have an some error boundary \(5\), then it's expressed like :\[ |v-65| \leq 5 \]Because this is the same as: \[ -5 \leq v-65 \leq 5 \]And adding \(65\) to both sides: \[ 65-5 \leq v \leq 65+5 \]So her actual speed is in between 5 less than 65 and 5 more than 65